# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ 5.6 What the iteration yields
Successor. Assume $M_{\alpha}\models\mathsf{SVC}(S)$ and let $X\in M_{\alpha+1}$ be nonempty. Choose a hereditarily symmetric $\mathbb{P}_{\alpha+1}$-name $\dot{X}\in\operatorname{HS}_{\alpha+1}$ (so $\dot{X}\in M_{\alpha}$) with $X=\dot{X}^{G_{\alpha+1}}$. By $\mathsf{SVC}(S)$ in $M_{\alpha}$, there exist an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow\dot{X}$ in $M_{\alpha}$.
Fix $p_{0}\in\mathbb{P}_{\alpha+1}$ and a name $\dot{x}_{0}$ such that $p_{0}\Vdash\dot{x}_{0}\in\dot{X}$. Define a $\mathbb{P}_{\alpha+1}$-name $\dot{h}$ for a function $S\times\eta\to\dot{X}$ by: for $s\in S$ and $\xi<\eta$, write $e(s,\xi)=\langle\dot{y},p\rangle$ and set
$\dot{h}(s,\xi)=\begin{cases}\dot{y}&\text{if }p\in\dot{G}_{\alpha+1},\\ \dot{x}_{0}&\text{otherwise.}\end{cases}$
Let $h:=\dot{h}^{G_{\alpha+1}}$. Then $h:S\times\eta\to X$, and $h$ is surjective by construction.
Moreover, $\dot{h}\in\operatorname{HS}_{\alpha+1}$: it is definable from parameters in $M_{\alpha}$, and any automorphism fixing $\dot{X}$ (hence lying in $\tilde{\mathscr{F}}^{*}_{\alpha+1}$) fixes $\dot{h}$ as well. Thus $h\in M_{\alpha+1}$, so $M_{\alpha+1}\models\mathsf{SVC}(S)$.
Limit. Let $\lambda$ be limit and assume $M_{\beta}\models\mathsf{SVC}(S)$ for all $\beta<\lambda$. Let $X\in M_{\lambda}$ be nonempty. By API item (e), $X$ has a hereditarily symmetric name with countable stage support bounded in $\lambda$, so $X\in M_{\beta}$ for some $\beta<\lambda$. By the inductive hypothesis, $M_{\beta}\models\mathsf{SVC}(S)$, hence $M_{\lambda}\models\mathsf{SVC}(S)$ as well.
@@
Proposition 5.71 (Final $\mathsf{AC}_{\mathsf{WO}}$).
$\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$.
@@
Work in the final symmetric model $\mathcal{M}$. By Lemma 5.70, we have $\mathcal{M}\models\mathsf{SVC}(S)$. We show that $\mathcal{M}\models\mathsf{AC}_{<\aleph^{*}(S)^{M}}$; by the Ryan–Smith equivalences this implies $\mathsf{AC}_{\mathsf{WO}}$.
Let $\lambda<\aleph^{*}(S)^{M}$ be an ordinal and let $f:Y\twoheadrightarrow\lambda$ be a surjection in $\mathcal{M}$. By $\mathsf{SVC}(S)$, fix an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow Y$ in $\mathcal{M}$. Let $g:=f\circ e:S\times\eta\twoheadrightarrow\lambda$.
Let $\dot{g}$ be a hereditarily symmetric name such that $g=\dot{g}^{\,G_{\infty}}$. Let $I\subseteq\mathrm{Ord}$ be the countable stage support of $\dot{g}$ (API item (e)), and let $\alpha:=\sup(I)+1$. Then $\dot{g}$ is a $\mathbb{P}_{\alpha}$-name, and $g_{\alpha}:=\dot{g}^{\,G_{\alpha}}\in M_{\alpha}$. By the bookkeeping (Definition 5.59), there is some successor stage $\beta>\alpha$ at which the orbit package $\mathbb{R}_{[g_{\alpha}]}$ is scheduled. Let $\dot{r}_{g_{\alpha}}$ be the canonical $\mathbb{R}_{[g_{\alpha}]}$-name for the generic right inverse in the $g_{\alpha}$-coordinate. By Lemma 5.46, the canonical $\mathbb{R}_{[g_{\alpha}]}$-name $\dot{r}_{g_{\alpha}}$ is hereditarily symmetric for the package symmetric system in $M_{\beta}$. By Lemma 5.47, its canonical $\mathbb{P}_{\beta+1}$-lift $\dot{r}_{g_{\alpha}}^{*}$ is hereditarily symmetric at stage $\beta+1$, hence $r_{g_{\alpha}}:=\dot{r}_{g_{\alpha}}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$, and $r_{g_{\alpha}}$ is a right inverse of $g$ in $\mathcal{M}$.
Now define $t:=e\circ r_{g_{\alpha}}:\lambda\to Y$. Then for all $\xi<\lambda$,
$f\bigl(t(\xi)\bigr)=f\bigl(e(r_{g_{\alpha}}(\xi))\bigr)=g\bigl(r_{g_{\alpha}}(\xi)\bigr)=\xi,$
so $t$ is a right inverse to $f$. Thus every surjection $Y\twoheadrightarrow\lambda$ with $\lambda<\aleph^{*}(S)^{M}$ splits in $\mathcal{M}$, and therefore $\mathcal{M}\models\mathsf{AC}_{<\aleph^{*}(S)^{M}}$.
Finally, since $\mathcal{M}\models\mathsf{SVC}(S)$, the Ryan–Smith theorem (see Remark 4.11) yields $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. ∎
Theorem 5.72 (Final model).
There is a transitive model $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
@@
By Theorem 5.56, $\mathcal{M}\models\mathsf{ZF}$, and by Proposition 5.65, $\mathcal{M}\models\mathsf{DC}$.
By Proposition 5.62, $\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$, hence $\mathcal{M}\models\mathsf{PP}\!\restriction T$. By Proposition 5.71, $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. By Lemma 5.70, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ for $T=\mathcal{P}(S)$, so the hypotheses of the Ryan–Smith localization theorem (Theorem 4.5, i.e., [17, Prop. 3.17] applied with Ryan–Smith’s parameter $S$ taken to be our $T$) hold in $\mathcal{M}$. Therefore $\mathcal{M}\models\mathsf{PP}$.
Finally, Proposition 5.64 gives $\mathcal{M}\models\neg\mathsf{AC}$. ∎

### v4

@@ 5.6 What the iteration yields
Successor. Assume $M_{\alpha}\models\mathsf{SVC}(S)$, and let $X\in M_{\alpha+1}$ be nonempty. Choose $\dot{X}\in\operatorname{HS}_{\alpha+1}$ with $X=\dot{X}^{G_{\alpha+1}}$. Since $\dot{X}\in M_{\alpha}$ (HS-names are formed in the stage ground), we may apply $\mathsf{SVC}(S)$ in $M_{\alpha}$ to the set $\dot{X}$ and obtain an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow\dot{X}$ in $M_{\alpha}$.
Fix $p_{0}\in\mathbb{P}_{\alpha+1}$ and a name $\dot{x}_{0}$ with $p_{0}\Vdash\dot{x}_{0}\in\dot{X}$. Work in the Boolean completion $\mathbb{B}_{\alpha+1}$ of $\mathbb{P}_{\alpha+1}$ in the background universe. For each $u=(s,\xi)\in S\times\eta$, write $e(u)=\langle\dot{y}_{u},p_{u}\rangle$. By the mixing lemma in $\mathbb{B}_{\alpha+1}$, there is a $\mathbb{P}_{\alpha+1}$-name $\dot{z}_{u}$ such that
$1\Vdash\dot{z}_{u}=\begin{cases}\dot{y}_{u}&\text{if }p_{u}\in\dot{G}_{\alpha+1},\\ \dot{x}_{0}&\text{otherwise.}\end{cases}$
Let $\dot{h}$ be the canonical name for the function $u\mapsto\dot{z}_{u}$ (so $\dot{h}$ is a $\mathbb{P}_{\alpha+1}$-name for a function $S\times\eta\to\dot{X}^{G_{\alpha+1}}$). Then $h:=\dot{h}^{G_{\alpha+1}}$ is a function $S\times\eta\to X$, and it is surjective: if $x\in X$, pick $\langle\dot{y},p\rangle\in\dot{X}$ with $p\in G_{\alpha+1}$ and $\dot{y}^{G_{\alpha+1}}=x$, and then pick $u$ with $e(u)=\langle\dot{y},p\rangle$; for that $u$, we have $h(u)=x$.
Moreover, $\dot{h}\in\operatorname{HS}_{\alpha+1}$: it is built in the stage ground $M_{\alpha}$ from $e$ and $\dot{X}$, and any automorphism in $\mathscr{G}_{\alpha+1}$ fixing $\dot{X}$ fixes the resulting mixed values $\dot{z}_{u}$ and hence $\dot{h}$. Therefore $h\in M_{\alpha+1}$, so $M_{\alpha+1}\models\mathsf{SVC}(S)$.
Limit. Let $\lambda$ be limit and assume $M_{\beta}\models\mathsf{SVC}(S)$ for all $\beta<\lambda$. Let $X\in M_{\lambda}$ be nonempty, and choose $\dot{X}\in\operatorname{HS}_{\lambda}$ with $X=\dot{X}^{G_{\lambda}}$. By the countable-support nature of the iteration and the HS-support analysis ((cf. the countable-support analysis of HS-names in Section 3 and the preceding discussion of supports at limits), $\dot{X}$ depends on only countably many stages. Let $\langle\lambda_{n}:n<\omega\rangle$ be an increasing sequence of ordinals below $\lambda$ capturing those stages. For each $n$, let $X_{n}$ be the interpretation of the corresponding stage-truncation of $\dot{X}$; then $X_{n}\in M_{\lambda_{n}}$ and $X=\bigcup_{n<\omega}X_{n}$.
By the inductive hypothesis, each $M_{\lambda_{n}}\models\mathsf{SVC}(S)$, so for each $n$ there is an ordinal $\eta_{n}$ and a surjection $e_{n}:S\times\eta_{n}\twoheadrightarrow X_{n}$ in $M_{\lambda_{n}}\subseteq M_{\lambda}$. Since $M_{\lambda}\models\mathsf{DC}$ (by Remark 2.4(d) and Lemma 5.54(iv)), $M_{\lambda}$ satisfies $\mathsf{AC}_{\omega}$, so we may choose such witnesses $\langle\eta_{n},e_{n}:n<\omega\rangle$ in $M_{\lambda}$. Let $\theta:=\sup_{n}\eta_{n}$ and define $e:S\times(\omega\cdot\theta)\twoheadrightarrow X$ by sending $(s,\omega\cdot\xi+n)$ to $e_{n}(s,\xi)$ when $\xi<\eta_{n}$, and to a fixed element of $X$ otherwise. Then $e\in M_{\lambda}$ and $e$ is surjective onto $X$. Hence $M_{\lambda}\models\mathsf{SVC}(S)$.
@@
Corollary 5.70.
Assume that $M_{\lambda}\models\mathsf{SVC}(S)$ for all $\lambda\in\mathrm{Ord}$. Then the final class-length model $\mathcal{M}$ satisfies $\mathsf{SVC}(S)$. Consequently, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ (for the relevant parameter $T$).
@@
Let $x\in\mathcal{M}$. Choose a set-sized HS name $\dot{x}\in\operatorname{HS}_{\mathrm{Ord}}$ with $x=\dot{x}^{G_{\infty}}$. By Lemma 5.64, letting $\alpha:=\sup(I(\dot{x}))+1$, we have $x=\dot{x}^{G_{\alpha}}\in M_{\alpha}$. Since $M_{\alpha}\models\mathsf{SVC}(S)$ (by Lemma 5.69), there is an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow x$ in $M_{\alpha}$, hence also in $\mathcal{M}$. Thus $\mathcal{M}\models\mathsf{SVC}(S)$. The final statement follows from Fact 4.3. ∎
Remark 5.71 (Structural characterization of $\mathsf{SVC}$ (context only)).
Blass [2] proved that $\mathsf{SVC}$ is equivalent to: $\mathsf{AC}$ is forceable over the model by a set forcing. We do not use deeper structural characterizations here; see, e.g., work in set-theoretic geology such as Usuba [19] for related perspectives.
In particular, Corollary 5.70 implies that $\mathsf{AC}$ is set-forceable over $\mathcal{M}$.
Proposition 5.72 ($\mathsf{AC}_{\mathsf{WO}}$ in the final model).
The final symmetric model $\mathcal{M}$ satisfies $\mathsf{AC}_{\mathsf{WO}}$.
@@
Let $f:S\times\eta\twoheadrightarrow\lambda$ be a surjection in $\mathcal{M}$, where $\eta,\lambda\in\mathrm{Ord}$. We show $f$ has a right inverse in $\mathcal{M}$.
Choose a set-sized hereditarily symmetric $\mathbb{P}_{\mathrm{Ord}}$-name $\dot{f}\in\operatorname{HS}_{\mathrm{Ord}}$ with $f=\dot{f}^{G_{\infty}}$. By Lemma 5.64, there is a set ordinal $\alpha$ such that $f\in M_{\alpha}$. If $f$ already has a right inverse in $M_{\alpha}$, we are done.
Otherwise, $f$ is eligible for scheduling by the $\mathsf{AC}_{\mathsf{WO}}$-bookkeeping clause in Definition 5.61. Hence by the guarantee clause in that definition, there exists a stage $\beta\geq\alpha$ such that either $f$ already has a right inverse in $M_{\beta}$, or else the successor iterand at stage $\beta$ includes the orbit $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{[f]}$. Assume the latter case and fix such $\beta$.
Let $\dot{r}_{f}$ be the canonical $\mathbb{R}_{[f]}$-name for the $f$-coordinate generic right inverse. In the full package extension, this coordinate generic yields a right inverse for $f$ (cf. Proposition 4.27). By Lemma 5.48, $\dot{r}_{f}$ is hereditarily symmetric for $(\mathbb{R}_{[f]},\mathscr{G}_{[f]},\mathscr{F}_{[f]})$ over $M_{\beta}$. By Lemma 5.49, the canonical $\mathbb{P}_{\beta+1}$-lift $\dot{r}_{f}^{*}$ is hereditarily symmetric at stage $\beta+1$. Therefore
$r_{f}:=\dot{r}_{f}^{*\,G_{\beta+1}}\in M_{\beta+1}\subseteq\mathcal{M}$
is a right inverse for $f$ in $\mathcal{M}$. Hence $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. ∎
Theorem 5.73 (Main theorem).
Let $V\models\mathsf{ZFC}$. In the class forcing extension $V[G]$ obtained by the $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen-seed symmetric model $\mathcal{N}$ described above, the final symmetric model $\mathcal{M}$ satisfies
$\mathsf{ZF}\ +\ \mathsf{DC}\ +\ \mathsf{PP}\ +\ \mathsf{AC}_{\mathsf{WO}}\ +\ \neg\mathsf{AC}.$
Proof.
@@
$\mathsf{ZF}$ holds in $\mathcal{M}$ by Theorem 5.58. $\mathsf{DC}$ holds by Proposition 5.68. $\mathsf{PP}$ holds by Theorem 4.5 together with Propositions 5.65 and 5.72. $\mathsf{AC}_{\mathsf{WO}}$ holds by Proposition 5.72. Finally, $\neg\mathsf{AC}$ holds by Proposition 5.67. ∎
Corollary 5.74 (Ordering Principle).
The final model $\mathcal{M}$ satisfies the Ordering Principle: every set admits a linear order.
Proof.
Work in $\mathcal{M}$. Fix any set $X$. By Corollary 5.70, $\mathcal{M}\models\mathsf{SVC}(S)$, so choose an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow X$.
By Theorem 4.5 together with Propositions 5.65 and 5.72, $\mathcal{M}\models\mathsf{PP}$. Hence there is an injection $i:X\hookrightarrow S\times\eta$.
Pulling back the lexicographic order on $S\times\eta$ along $i$ yields a linear order on $X$. ∎
Remark 5.75 (Kinna–Wagner principles).
For $n\geq 1$, the Kinna–Wagner principle $KWP_{n}$ states: for every set $X$ there exists an ordinal $\theta$ such that $X\hookrightarrow\mathcal{P}^{n}(\theta)$ (where $\mathcal{P}^{n}$ is the $n$-fold iterated powerset).
(1) $KWP_{2}$ from $\mathsf{SVC}(S)$. Work in $\mathcal{M}$ and fix a set $X$. By Corollary 5.70, choose an ordinal $\eta$ and a surjection $e:S\times\eta\twoheadrightarrow X$. The map
$x\ \longmapsto\ e^{-1}(\{x\})$
is an injection of $X$ into $\mathcal{P}(S\times\eta)$. Since $S=A^{\omega}$ injects into $2^{\omega}$, we have $S\times\eta\hookrightarrow 2^{\omega}\times\eta$. Identifying $2^{\omega}$ with $\mathcal{P}(\omega)$, define an injection
$\mathcal{P}(\omega)\times\eta\ \hookrightarrow\ \mathcal{P}(\omega\times\eta),\qquad(r,\xi)\longmapsto\{(n,\xi):n\in r\}.$
Thus $S\times\eta\hookrightarrow\mathcal{P}(\omega\times\eta)$, so $\mathcal{P}(S\times\eta)\hookrightarrow\mathcal{P}(\mathcal{P}(\omega\times\eta))=\mathcal{P}^{2}(\omega\times\eta)$. Because $\omega\times\eta$ is well-orderable of type $\omega\cdot\eta$, we have $\mathcal{P}^{2}(\omega\times\eta)\cong\mathcal{P}^{2}(\omega\cdot\eta)$, and hence $X\hookrightarrow\mathcal{P}^{2}(\theta)$ for $\theta=\omega\cdot\eta$. This proves $KWP_{2}$ in $\mathcal{M}$.
(2) $KWP_{1}$ using $\mathsf{PP}$. By Theorem 4.5 together with Propositions 5.65 and 5.72, $\mathcal{M}\models\mathsf{PP}$. Apply $\mathsf{PP}$ to the surjection $e:S\times\eta\twoheadrightarrow X$ to obtain an injection $i:X\hookrightarrow S\times\eta$. Composing with the injection $S\times\eta\hookrightarrow\mathcal{P}(\omega\times\eta)\cong\mathcal{P}(\omega\cdot\eta)$ from (1), we obtain $X\hookrightarrow\mathcal{P}(\theta)$ for $\theta=\omega\cdot\eta$. Thus $KWP_{1}$ holds in $\mathcal{M}$.
6 Conclusion
Starting from a ground model $V\models\mathsf{ZFC}$, we define an $\mathrm{Ord}$-length countable-support symmetric iteration over the Cohen symmetric seed model $\mathcal{N}$ and let $\mathcal{M}$ be the resulting final symmetric model. The iteration is arranged so that $\mathcal{M}$ satisfies $\mathsf{DC}$ (Proposition 5.68), satisfies $\mathsf{AC}_{\mathsf{WO}}$ (Proposition 5.72), and still contains the non-well-orderable Cohen set $A$ (Proposition 5.67). We force the localized splitting principle $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ for the fixed parameter $T$ (Proposition 5.65) and verify $\mathsf{SVC}^{+}(T)$ in $\mathcal{M}$ via persistence of $\mathsf{SVC}(S)$ (Lemma 5.69 and Corollary 5.70). Ryan–Smith localization (Theorem 4.5) then yields $\mathcal{M}\models\mathsf{PP}$, completing the proof that $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\mathsf{AC}_{\mathsf{WO}}+\neg\mathsf{AC}$ (Theorem 5.73). As recorded in the final section, this also implies the Ordering Principle (Corollary 5.74) and the stated Kinna–Wagner consequences (Remark 5.75).